Hayhayz (hayhayz):

what are the similarities and differences of linear and non linear systems of equations? @dan815

OpenStudy (dan815):

x+y=2 2x+3y=1 ------------------- x^2+y^2=2 2x^3+3y^4=1 ------------------------ The general methods such as substitution and elimination still apply to both equations. However for a linear set of equations it is easier and more clear to see the form of the solution. For example 2 lines can intersect at a point, or overlap or be parallel. For 2 non linear equations, a multiple of things can occur, and the number of solutions and the form of solutions can greatly vary like in the case of a multiple peaked slope intersection through the circumference of a circle. Also there can be complex solutions, and it is not always possible to solve non linear equations analytically.

Hayhayz (hayhayz):

ty

OpenStudy (dan815):